OpenStudy (anonymous):

If 6 is subtracted from the third of three consecutive odd integers and the result is multiplied by 2, the answer is 31 less than the sum of the first and twice the second of the integers. Find the integers. the first integer is

OpenStudy (jack1):

(x+2y)-31 = 2(z-6)

OpenStudy (jack1):

consecutive....how ie next odd number ascending from 0 ... so y = x+2 and z = y + 2 ...?

OpenStudy (jack1):

@tryder ?

OpenStudy (jack1):

ok, so lets assume that it is consecutive in they're ascending odd numbers so intergers are x, y, z y = x+2 z = y + 2 so (x+2y)-31 = 2(z-6) (x+2[x+2])-31 = 2([y+2]-6) (x+2[x+2])-31 = 2([{x+2}+2]-6) (x+2x+4-31 = 2(x+2+2-6) 3x-27 = 2x+4+4-12 3x-27 = 2x-4 x = 31

OpenStudy (jack1):

so x = 31, y = 33, z = 35...?

OpenStudy (yttrium):

(x+1)+2(x+3)-31 = 2(x+5-6) x+1+2x+6-31 = 2x-2 x = 22 x+1 = 23 x+3 = 25 x+5 = 27

OpenStudy (yttrium):

To check: 23+2(25)-31 = 2(27-6) 73-31 = 2(21) 42 = 42 Therefore, correct is 23, 25, 27. :)

OpenStudy (jack1):

yep, hes right... where's my math wrong tho...?

OpenStudy (jack1):

no, found it, last line was: 3x-27 = 2x+4+4-12 3x-27 = 2x-4 x = 31 should be 3x-27 = 2x+4+4-12 3x-27 = 2x-4 x = -23

OpenStudy (jack1):

duh, typo x = 23

OpenStudy (yttrium):

it's no typo, mental calculation error it is

OpenStudy (jack1):

aye, cheers @Yttrium slaters

OpenStudy (whpalmer4):

can be done a bit more simply: let numbers be \(x,x+2,x+4\) and just verify that result is odd \[(x+4-6)*2+31 = x+2(x+2)\]\[2x+27=x+2x+6\]\[27-6=x+2x-2x\]\[x=23\]which is odd, as required, giving us our answer of \(23,25,27\) @jack1 your error was in solving \[3x-27=2x-4\]You didn't change the sign on the -27 when you shifted it to the other side of the = sign. \[3x-27=2x-4\]subtract \(2x\) from each side giving\[x-27=-4\]add 27 to each side giving\[x=-4+27 = 23\]

OpenStudy (jack1):

to true, cheers man @whpalmer4

OpenStudy (yttrium):

as i say, mental calculation error it is